### v1

# Image Corruption-Inspired Membership Inference Attacks against Large Vision-Language Models

Abstract
Large vision-language models (LVLMs) have demonstrated outstanding performance in many downstream tasks. However, LVLMs are trained on large-scale datasets, which can pose privacy risks if training images contain sensitive information. Therefore, it is important to detect whether an image is used to train the LVLM. Recent studies have investigated membership inference attacks (MIAs) against LVLMs, including detecting image-text pairs and single-modality content. In this work, we focus on detecting whether a target image is used to train the target LVLM. We design simple yet effective Image Corruption-Inspired Membership Inference Attacks (ICIMIA) against LLVLMs, which are inspired by LVLM’s different sensitivity to image corruption for member and non-member images. We first perform an MIA method under the white-box setting, where we can obtain the embeddings of the image through the vision part of the target LVLM. The attacks are based on the embedding similarity between the image and its corrupted version. We further explore a more practical scenario where we have no knowledge about target LVLMs and we can only query the target LVLMs with an image and a question. We then conduct the attack by utilizing the output text embeddings’ similarity. Experiments on existing datasets validate the effectiveness of our proposed attack methods under those two different settings.
Image Corruption-Inspired Membership Inference Attacks against
Large Vision-Language Models
Zongyu Wu, Minhua Lin, Zhiwei Zhang, Fali Wang, Xianren Zhang, Xiang Zhang, Suhang Wang The Pennsylvania State University {zongyuwu,mfl5681,zbz5349,fqw5095,xzz5508,xzz89,szw494}@psu.edu
1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2024a, b; Bai et al., 2023b; Achiam et al., 2023), which can generate text outputs based on visual and/or textual input, are attracting increasing attention. Many LVLMs have been developed, which have shown great performance on various tasks (Dong et al., 2024; Xu et al., 2024; Wu et al., 2025; Yue et al., 2024; Li et al., 2024a; Bucciarelli et al., 2024) such as biomedical question answering (Li et al., 2023a).
Figure 1: An example of some potential risks in LVLMs.
Despite their great success, LVLMs have also raised critical concerns about privacy and copyright issues as shown in Figure 1. LVLMs are usually trained on large-scale datasets (Ng et al., 2020; Changpinyo et al., 2021; Byeon et al., 2022). The training data might contain sensitive information, such as unauthorized medical data and copyrighted content. Previous works have shown that neural networks, especially large models, can memorize training data (Song et al., 2017; Carlini et al., 2019, 2022b). Thus, the memorization phenomena might also happen in LVLMs and inadvertently cause training data leakage (Li et al., 2024b), which could cause substantial loss to data owners. Hence, knowing whether one’s data is used to train an LVLM is important for privacy and copyright protection.
Membership inference attacks (MIAs), which aim to determine whether a given sample is used to train a model (Hu et al., 2022a; Shokri et al., 2017), are critical for ensuring data integrity (Oren et al., 2023; Duan et al., 2024). Generally, models tend to overfit to the training data, resulting in higher prediction confidence for member data (data used for training) than non-member data. Many traditional MIA methods adopt such nuance difference in model prediction to differentiate members and non-members.
Recently, some works (Ko et al., 2023; Li et al., 2024b; Hu et al., 2025) have explored the MIA on vision-language models, from CLIP (Radford et al., 2021) to LVLMs (Zhang et al., 2023). Ko et al. (2023) focus on determining whether an image-text pair exists in the training data of CLIP. However, detecting whether an image in the training data is more practical than detecting the entire image-text pair as the image owner might only have the image without text while the text description might be labeled by the model trainer (Li et al., 2024b).
Therefore, we study the problem of single-image based MIA against LVLMs. The work in this direction is rather limited Li et al. (2024b). Li et al. (2024b) also conduct MIA on a single modality (Image or textual description) against LVLMs by using the output logits of LVLM. Instead of using output logits, we propose a new perspective, i.e., adopting image embedding from the visual part of LVLMs. Our intuition is: as the model has seen member images, it should be able to give robust image embedding of a member image even if some details of the image are missing. In other words, the image embedding of a member is more robust to image corruption than that of non-member images, which is verified by our preliminary experiment in Section 3 (see Fig. 3). Based on this observation, we propose a novel MIA algorithm under the white-box setting, where we can obtain the image embedding from LVLMs. Given an image, we corrupt the image and utilize the image embedding similarity between the raw image and the corrupted version to determine if an image is a member. A higher similarity means the image is more likely to be a member.
However, for many closed-source LVLMs, we cannot obtain image embeddings. To address this issue, we extend the similarity to the output text level. Our assumption is: robust image embedding of member image will result in robust text generation under perturbation, which is also verified in Figure 4. Based on this observation, we extend our framework to black-box setting, where we can only obtain output texts from LVLMs. Given a target image, we corrupt the image and compare the generated text similarity between the raw image and its corrupted version. A larger text similarity means the image is more likely to be a member.
Our main contributions are: (i) In this work, we investigate two membership inference attack scenarios targeting LVLMs. For each setting, we propose one simple yet strong attack method that leverages the model’s robustness to image corruption on its training images; (ii) Extensive experiments on existing datasets show the effectiveness of the proposed method.
2 Related Works
2.1 Large Vision-Language Models
Large Vision-Language Models (Bai et al., 2025; Chen et al., 2024a, b; Liu et al., 2024b, a; Chen et al., 2024c; Tong et al., 2024; Zhu et al., 2025; Young et al., 2024; Zhu et al., 2023; Chen et al., 2023; Li et al., 2024c; Du et al., 2025), also known as multimodal large language models (Fu et al., 2024) are developing rapidly due to the success of language models (Wang et al., 2024; Zhao et al., 2023; Chiang et al., 2023; Touvron et al., 2023a, b; Grattafiori et al., 2024; Yang et al., 2025, 2024; Jiang et al., 2023; Team, 2023; Penedo et al., 2023) such as Qwen (Bai et al., 2023a; Yang et al., 2024) and LLaMA series (Touvron et al., 2023a, b). These models, like LLaVA 1.5 (Liu et al., 2024c), can generate textual outputs given textual questions and images.
2.2 Membership Inference Attack
Membership Inference Attack (MIA) (Shokri et al., 2017; Salem et al., 2018; Sablayrolles et al., 2019; Li et al., 2021; Hu et al., 2022a; Nasr et al., 2019; Leino and Fredrikson, 2020; Rezaei and Liu, 2021) tries to determine whether a given data sample was used to train a machine learning model (Shokri et al., 2017).
One main category of MIA methods is metric-based (Hu et al., 2022a; Li et al., 2024b; Sablayrolles et al., 2019; Choquette-Choo et al., 2021; Li et al., 2021), which use some well-designed metrics, such as the entropy of output logits, to determine the membership status of a given sample. Another main category is methods based on shadow training (Shokri et al., 2017), which trains some shadow models to simulate the target model and then conducts MIAs based on these shadow models. Many later works have started to explore the MIA on LLMs (Mattern et al., 2023; Mireshghallah et al., 2022; Ren et al., 2024; Shi et al., 2024; Zhang et al., 2024) such as Min-K(Shi et al., 2024) and Min-K%++ (Zhang et al., 2024).
With the development of multimodal learning, there are also some works exploring MIA on multimodal models (Hu et al., 2022b; Ko et al., 2023; Li et al., 2024b). We focus on vision-text here. Liu et al. (2021) studies MIAs on image encoder models such as CLIP vision encoder (Radford et al., 2021). It calculates the similarity scores between the augmented images’ embeddings and the scores are then used to train a classifier to infer the member status of an image. Ko et al. (2023) work on detecting whether an image-text pair is in the training data of CLIP models (Radford et al., 2021). Li et al. (2024b) investigates the single-modality MIA in LVLMs, which is a more practical scenario. They calculate the Rényi Entropy (Rényi, 1961) based on different slices of logits. Hu et al. (2025) found that member data and non-member data have different sensitivity to temperature. They then perform four attack methods under four different settings to detect whether a group of images is used in the training stage.
Our work is inherently different from existing work: We propose a new perspective from image embedding robustness and output text robustness of the member image under image corruption, which works for both white-box and black-box settings.
3 Preliminaries
In this section, we give the necessary background information and formulate the problems.
Figure 2: An illustration of the architecture of large vision-language models (Liu et al., 2024a).
3.1 Large Vision-Language Models
As shown in Fig. 2, an LVLM $\mathbf{M}_{\theta}$ usually consists of three parts: A vision encoder $f_{Vision}$, an LLM $f_{LLM}$, and a modality connection module $f_{Align}$. The vision encoder $f_{Vision}$, e.g., CLIP Radford et al. (2021), takes an image $\mathbf{x}$ as input and outputs the embeddings of $N$ image patches $\{\mathbf{z}_{1},...,\mathbf{z}_{N}\}$, where $\mathbf{z}_{i}\in\mathbf{R}^{d_{v}}$ is the $i$-th patch embedding with $d_{v}$ being the embedding dimension. The patch embeddings are then transformed into an embedding sequence $\mathbf{E}^{v}=\{\mathbf{e}_{1}^{v},...,\mathbf{e}_{N}^{v}\}$ which is in the embedding space of $f_{LLM}$ using $f_{Align}$. Each element in $\mathbf{E}^{v}$ can be represented as:
$\mathbf{e}_{i}^{v}=f_{Align}(\mathbf{z}_{i}),$ (1)
The text prompt $T_{in}$, composed of the instruction and the question, is encoded into an embedding sequence of tokens $\mathbf{E}^{t}=\{\mathbf{e}_{1}^{t},...,\mathbf{e}_{K}^{t}\}$, where $K$ is the number of tokens. Finally, the text embedding and the image embeddings are used as input to the LLM $f_{LLM}$ to get the text output as $T_{out}=f_{LLM}(\mathbf{E}^{v},\mathbf{E}^{t})$.
Figure 3: A histogram of similarity scores of corrupted images’ (Through Gaussian Blur) embeddings and original images’ embeddings for member and non-member data. Scores in this figure achieve an AUC of 0.881.
Figure 4: A histogram of similarity scores of textual output embeddings between corrupted images (Through Gaussian Blur) and the original images for member and non-member data. Scores in this figure achieve an AUC of 0.652.
3.2 Threat Model
Attacker’s Goal. Given a trained LVLM $\mathbf{M}_{\theta}$, the attacker’s goal is to determine if a specific image was in the training set of the target LVLM $\mathbf{M}_{\theta}$. The attackers only have the target image and do not require its corresponding ground-truth text description. The problem is formally defined as:
Definition 1 (Image Only MIA)
Given a trained LVLM $\mathbf{M}_{\theta}$ and a target image $\mathbf{x_{t}}$, the attacker aims to determine whether $\mathbf{x_{t}}$ was part of the training data for $\mathbf{M}_{\theta}$, i.e., $\text{Attack}(\mathbf{M}_{\theta},\mathbf{x_{t}},T_{in})\rightarrow\{0,1\}$, where $T_{in}$ is the input textual instruction. $\mathbf{x_{t}}$ is considered to be in the training set (i.e, member image) if the output is 1. Otherwise, $\mathbf{x}_{t}$ is a non-member.
Attacker’s Knowledge. In our work, we consider two practical settings, i.e., white-box and black-box. (i) White-box Setting. In this setting, the attacker can get the embedding of the image, i.e., the embedding through the model’s vision encoder and modality alignment module. This is practical for open-source LVLMs. (ii) Black-box Setting. In this setting, the attacker can only query the LVLM $\mathbf{M}_{\theta}$ with image input $\mathbf{x}$ and text input $T_{in}$ to get the text output $T_{out}$. The attacker has no knowledge of the target model. This setting is more realistic for commercial LVLMs.
3.3 Intuition for Performing MIA
As we are only given the image, how to fully utilize the image to determine if the image remains a challenging question. Inspired by Liu et al. (2021) that a CLIP vision encoder tends to overfit to its training data and member images show higher similarity for two augmented versions, we assume that the image embedding of a member image from the vision encoder and modality alignment module should be more robust to image corruption than that of non-member images. The intuition is, if the model has seen the image and memorized the image, then even if the image is corrupted, e.g., some details missing, the model can still recall the details and result in embedding similar to the original one. To verify our assumption, for each image, we first apply Gaussian Blur (Bradski, 2000) as the corruption method to obtain the corrupted image, where Gaussian Blur is a smoothing technique that reduces image detail and noise by averaging pixels with a weighted Gaussian kernel. We compute the similarity score between the embedding of the original image and that of its corrupted version. The chosen model is LLaVA-1.5-7B (Liu et al., 2024a), and the dataset is the Img_Flickr dataset constructed by Li et al. (2024b). The similarity scores of member images and non-member images are shown in Figure 3. We can observe that member images normally show a higher similarity score than non-member images, which aligns well with our assumption. The scores show a clear difference between the two groups, which suggests that the similarity score between the original image embedding and its corrupted version’s embedding can be used to perform an image membership inference attack. Thus, we can utilize the robustness of member image embedding to determine its membership.
However, in black-box setting, we are unable to obtain the image embedding. As image embedding of member image is robust, it might also result in robust text even if the image is corrupted. Thus, we assume that given the same text prompt (Instruction), an image seen during the LVLM’s training process will produce a text output that is more similar to the output generated from its corrupted version, compared to images not included in the training data. To verify our assumption, we use the same dataset, model, and corruption method as above. We compute the similarity score between the output text embeddings of the original and corrupted versions of each image, given the same prompt, separately. The results are shown in Figure 4. We observe that the discrepancy in similarity between member and non-member images still exists and can serve as a metric, although it is less apparent in Figure 3.
Figure 5: An illustration of the attack pipeline under two different settings.
4 Method
Based on observations that LVLMs are more robust to corruptions on members than non-members, we propose a novel framework, ICIMIA (Image Corruption-Inspired Membership Inference Attacks against Large Vision-Language Models). As shown in Figure 5, our method first produces a corrupted version $\mathbf{x}_{t}^{\prime}$ for the target image $\mathbf{x}_{t}$. Then both the $\mathbf{x}_{t}$ and $\mathbf{x}_{t}^{\prime}$ are utilized to get their image embeddings or corresponding output text embeddings. Finally, we calculate the image/text embedding similarity as a metric. The image is viewed as a member image if the similarity score is bigger than a certain threshold. Next, we introduce the details.
4.1 White-Box MIA via Image Similarity
Algorithm 1 Image Similarity-based MIA
1:Target image $\mathbf{x}_{t}$, threshold $\lambda$
2:Membership Prediction Result $\in\{0,1\}$
3:Obtain the embedding $\mathbf{E}^{v}$ of $\mathbf{x}_{t}$
4:Apply corruption on $\mathbf{x}_{t}$ to get image $\mathbf{x}_{t}^{\prime}$
5:Obtain the embedding $\mathbf{E}^{v^{\prime}}$ of $\mathbf{x}_{t}^{\prime}$
6:Compute similarity score $s_{img}$ via Eq. 2
7:if $s_{img}<\lambda$ then
8:     $\mathbf{x}_{t}$ is a viewed as a non-member image
9:else
10:     $\mathbf{x}_{t}$ is a viewed as a member image
11:end if
We first discuss the grey-box setting where we can obtain the embeddings of a given image through the vision encoder and modality alignment module.
Based on our observation in Section 3.3, the embeddings of trained images should be more robust to image corruption methods. To be specific, compared to images not used during training (Non-member data), the embeddings of the corrupted image and its original counterpart are normally more similar for images that were used in the training process (Member data). Therefore, we can use the similarity score as a metric to determine whether an image is used to train the LVLM.
The pipeline of our proposed framework is shown in the upper part of Figure 5. For a target image $\mathbf{x}_{t}$ that we wanna detect, we first apply some corruptions to get a corrupted image as $\mathbf{x}_{t}^{\prime}=\text{Corruption}(\mathbf{x}_{t})$. Various corruption techniques can be used, such as Gaussian blur (Using a Gaussian Kernel). Then, we get the image embeddings for both the original image and the corrupted one. Last, we measure how close each patch embedding pair is and calculate the mean value as
$s_{img}=\frac{1}{N}\sum_{i=1}^{N}\text{Sim}(\mathbf{e}_{i}^{v},\mathbf{e}_{i}^{v^{\prime}})$ (2)
where $\mathbf{e}_{i}^{v}$ is the embedding of the $i$-patch of image $\mathbf{x}_{t}$ obtained by Eq. 1 and $\mathbf{e}_{i}^{v^{\prime}}$ denotes the $i$-patch embedding of the corrupted image $\mathbf{x}_{t}^{\prime}$. $N$ is the number of patches. Sim is a similarity function. We use cosine similarity here. The larger the $s_{img}$ is, the more likely the image is viewed as a member image. The target image $\mathbf{x}_{t}$ is predicted as a member image if $s_{img}$ is bigger than a threshold $\lambda$.
Algorithm 1 summarizes the image similarity-based attack.
4.2 Black-box MIA via Text Similarity
For many commercial LVLMs, we are not able to obtain image embeddings. Thus, we study a more practical black-box setting where we know nothing about the target model but can only query the model with the target image and prompt to obtain the response text. Though we cannot obtain the image embedding, as the image embeddings of member images are robust to random corruption, correspondingly, the generated response text will be robust to the corruption, which is also verified in 3.3. This motivates us to query the target model using the original image and its corrupted version. Then we calculate pair-wise output text similarities.
Specifically, we feed the original target image $\mathbf{x}_{t}$ and a text prompt $T_{in}$ into the target LVLM and get an output $T_{out}$ as $T_{out}=f_{LVLM}(\mathbf{x}_{t},T_{in})$. Similarly, we get an output $T_{out}^{\prime}$ for the corrupted image $\mathbf{x}_{t}^{\prime}$ with the same input text prompt $T_{in}$. We then employ a text embedding model to get the embeddings of $T_{out}$ and $T_{out}^{\prime}$. With the text embedding, we calculate their similarity as
$s_{text}=\text{Sim}(Emb(T_{out}),Emb(T_{out}^{\prime}))$ (3)
where $Emb()$ denotes a text embedding model, such as OpenAI’s text-embedding-3-small model [^1] https://platform.openai.com/docs/models/text-embedding-3-small. Sim is a similarity function as in Eq. 2 which is a cosine similarity function. Similarly, $x_{t}$ is considered member data if $s_{text}>\lambda$. This method is summarized in Algorithm 2.
Algorithm 2 Text Embedding Similarity-based Attack
1:Target image $\mathbf{x}_{t}$, Prompt $T_{in}$, threshold $\lambda$
2:Membership Prediction Result $\in\{0,1\}$
3:Feed the target LVLM with the target image $\mathbf{x}_{t}$ and a text prompt $T_{in}$ and get output $T_{out}$
4:Apply corruption to get corrupted image $\mathbf{x}_{t}^{\prime}$
5:Feed the target LVLM with the corrupted target image $\mathbf{x}_{t}^{\prime}$ and the same text prompt $T_{in}$ and get output $T^{\prime}_{out}$
6:Get the text output embedding $Emb(T_{out})$ and $Emb(T_{out}^{\prime})$ of the original image and corrupted image
7:Calculate the output similarity using Eq. 3
8:if $s_{text}<\lambda$ then
9:     $\mathbf{x}_{t}$ is viewed as a non-member image
10:else
11:     $\mathbf{x}_{t}$ is viewed as a member image
12:end if
In this section, we evaluate our methods on existing datasets to answer the following research questions: (i) How well do our proposed methods perform in conducting membership inference attacks against large vision-language models? and (ii) What are the impacts of hyperparameters?
4.3 Experimental Setup
Evaluation Metric. Following previous work Li et al. (2024b), we use Area Under the Curve (AUC) and True Positive Rate at 5% False Positive Rate (TPR at 5% FPR) as the evaluation metrics. For both metrics, higher values indicate better MIA performance. The descriptions of these metrics can be found in Appendix C.
Datasets We test our method on two benchmark datasets, VL-MIA/Flickr and VL-MIA/Flickr-2k (Li et al., 2024b). The details of datasets can be found in Appendix A.
Computational Resources. We conduct all experiments on machines equipped with four NVIDIA RTX A6000 GPUs (48GB memory).
Selected Models. We evaluate our method on two models: LLaVA 1.5 7B, and LLaVA 1.5 13B (Liu et al., 2024a). The models are chosen as they are classical and the datasets are applicable to them.
Corruption Methods. We choose the following corruption methods: (i) Gaussian Blur: This is a technique that makes image soft and blurry using a Gaussian Kernel; (ii) Motion Blur: We apply a custom convolution kernel, mimicking the effect of horizontal motion-blur; and (iii) JPEG Compression: This is a method to compress the image to get the corrupted version. We apply CV2 (Bradski, 2000) to achieve Gaussian Blur and Motion Blur. We put some examples of different corruption techniques in Appendix B.
Baselines. We adopt representative and state-of-the-art MIA methods against LVLMs, including: (i) AugKL (Li et al., 2024b; Liu et al., 2021)Li et al. (2024b) extend the approach designed by Liu et al. (2021) to LVLMs. Specifically, Li et al. (2024b) quantify the difference between the original and augmented images (Such as Crop and Rotation) by computing the KL divergence between their distributions of logits; (ii) Max_Prob_Gap (Li et al., 2024b): It is the average value of the difference between the highest and second-highest token probabilities at each position. (iii) Min-K (Shi et al., 2024): Min-K is an MIA method designed for LLMs. It uses a ground-truth token and computes the lowest Kof its predicted probabilities. Li et al. (2024b) extends it to the LVLM domain; (iv) Perplexity: It is based on loss. Li et al. (2024b) analyze target perplexity to achive the attacks (Carlini et al., 2021); (v) MaxRényi-K (Li et al., 2024b): It selects the top Ktokens with the highest Rényi entropy. Then the value is the average value of these entropies; and (vi) Mod-Rényi (Li et al., 2024b): This is an extended version of MaxRényi-K which is designed for target-based scenarios.
5 Experiment
Table 1: Image MIA AUC of various baseline methods under Li et al. (2024b)’s pipeline and our proposed approach on VL-MIA/Flickr. For all the baselines, the term "img" refers to the logits segment associated with the image embedding, while "inst" represents the instruction segment. "desp" corresponds to the generated description’s logits segment, and "inst+desp" denotes the combination of the instruction and description segments.
Method LLaVA-1.5-7B LLaVA-1.5-13B
img inst desp inst+desp img inst desp inst+desp
Perplexity - 0.378 0.665 0.558 - 0.440 0.707 0.646
Min_20Prob - 0.374 0.672 0.370 - 0.454 0.684 0.433
ModRényi - 0.370 0.658 0.613 - 0.442 0.703 0.678
Max_Prob_Gap 0.579 0.605 0.644 0.645 0.565 0.501 0.656 0.652
Aug_KL 0.665 0.568 0.537 0.549 0.636 0.540 0.538 0.552
MaxRényi 0.702 0.726 0.709 0.743 0.647 0.682 0.728 0.738
Ours (Image_Sim, Gaussian Blur, Kernel Size 5) 0.881 0.878
Ours (Image_Sim, Motion Blur Kernel Size 5) 0.860 0.856
Ours (Image_Sim, JPEG Compression, Quality = 5) 0.682 0.681
Table 2: TPR at 5% FPR of various baseline methods under Li et al. (2024b)’s pipeline and our proposed approach on VL-MIA/Flickr. The column ’img’, ’inst’, ’desp’, and ’inst+desp’ has the same meaning as the previous table.
Method LLaVA-1.5-7B LLaVA-1.5-13B
img inst desp inst+desp img inst desp inst+desp
Perplexity - 0.007 0.137 0.067 - 0.047 0.227 0.127
Min_20Prob - 0.007 0.127 0.003 - 0.067 0.163 0.053
ModRényi - 0.003 0.113 0.113 - 0.060 0.203 0.147
Max_Prob_Gap 0.050 0.083 0.163 0.163 0.050 0.107 0.163 0.160
Aug_KL 0.080 0.073 0.060 0.043 0.133 0.070 0.050 0.060
MaxRényi 0.100 0.210 0.163 0.127 0.077 0.073 0.213 0.183
Ours (Image_Sim, Gaussian Blur, Kernel Size 5) 0.333 0.323
Ours (Image_Sim, Motion Blur Kernel Size 5) 0.363 0.297
Ours (Image_Sim, JPEG Compression, Quality = 5) 0.057 0.060
Implementation details. All these baselines need the knowledge of the target model’s tokenizer and the output logits. The implementations of all baselines are based on Li et al. (2024b). Following the recommendation of Shi et al. (2024) and Li et al. (2024b), we set $K=20$ for the Min-K method. $K$ is set to 0, 10, and 100 for MaxRényi-$K\%$ and $\alpha$ is chosen over 0.5, 1, and 2 for both ModRényi and MaxRényi-$K\%$, as in Li et al. (2024b). We report the highest AUC for each method and provide the TPR at 5% FPR using the hyperparameter combination that achieves this highest AUC. For the Image Similarity-based attack, we fix the Kernel Size of the Gaussian blur and the Motion blur as 5. For JPEG compression, the image quality is set to 5 (lower values indicate stronger compression). For the Text Similarity-based attack, following Li et al. (2024b), we use “Describe this image concisely” (Li et al., 2024b) as the prompt and the max generation token amount is 32. For simplicity, we only use Gaussian blur for Text Similarity-based attack and the kernel size is set to 45.
Figure 6: Text similarity-based membership inference attack on VL-MIA/Flickr. Image corruption method is Gaussian Blur and the kernel size is 45.
Figure 7: Analysis on Image Corruption Hyperparameters.
5.1 White-Box MIA Performance
The results on LLaVA 1.5 series are shown in Table 1 and Table 2. Results on VL-MIA/Flickr 2K are in Appendix D. Note that our methods have different knowledge of the models compared to the baseline methods. We can obtain the image embeddings while they can obtain the output text’s logits. We can observe that similarity-based attacks using Gaussian Blur and Motion Blur can achieve an AUC higher than 0.8 for both LLaVA 1.5-7B and LLaVA-1.5-13B, which is much higher than all the baselines. In comparison, the highest AUC among the baseline methods is 0.743 on LLaVA-1.5-7B and 0.738 on LLaVA-1.5-13B. Similar results can be found in Table 2, our methods largely outperform all the baselines in terms of TPR at 5% FPR where the best baseline performance is 0.210 on LLaVA-1.5-7B and 0.227 on LLaVA-1.5-13B.
It shows that using information from the visual side can better facilitate image MIAs compared with using information from different logit slices. The results also suggest that the image corruption method matters. For example, the performance using JPEG compression is worse than that using Gaussian Blur and Motion Blur.
Figure 8: Kernel Size’s Impact on Text Similarity-based Attack.
5.2 Black-box MIA performance
The results of our attack method under the black-box setting, which is based on output texts’ embeddings, are shown in Figure 6. The results demonstrate the effectiveness of our method. For example, it achieves an AUC of 0.652 and a TPR@5%FPR of 0.137 on LLaVA-1.5-7B, which outperforms many baselines that require using output logits, even though our approach only queries the target model.
5.3 Hyperparameter Analysis
In this subsection, we conduct experiments to observe the impacts of hyperparameters. For the image similarity-based attack, we vary the kernel size for Gaussian Blur and Motion Blur using values of 3, 5, 7, 15, 25, 35, and 45. The image quality for JPEG compression is selected across 1, 5, 10, 20, 45, and 70. The selected model is LLaVA 1.5 7B. The results are shown in Figure 7. We can observe that smaller kernel sizes generally contribute to better performance in terms of both AUC and TPR@5%FPR. Larger kernel sizes will bring stronger corruption. Therefore, for blur-based corruption methods, it suggests that it would be better to use a smaller corruption level for such methods. Although the performance drops a lot with very large kernel sizes, it still outperforms most baselines that rely on output logits. For the third sub-figure in Figure 7, we can see that a smaller JPEG quality (Stronger compression) generally leads to higher AUC.
For the text similarity-based attack, the kernel size for Gaussian Blur is varied across 15, 25, 35, 45 and 55. The results are shown in Figure 8. Unlike image similarity-based attacks, text similarity-based attacks generally perform better with larger kernel sizes. One possible reason is that if the kernel size is too small, small changes in the image embeddings lead to even smaller changes in the textual output. This makes the difference between member and non-member images less clear.
Table 3: Image MIA AUC and TPR at 5% FPR of our proposed approach on VL-MIA/Flickr. The embeddings are replaced with the ones obtained directly from the CLIP vision encoder without the alignment module.
Method AUC TPR@5%FPR
7B 13B 7B 13B
Ours (Gaussian Blur) 0.881 0.878 0.333 0.323
Ours (Gaussian Blur, CLIP) 0.885 0.413
[…]
Ours (Motion Blur, CLIP) 0.858 0.393
[…]
Ours (JPEG Compression, CLIP) 0.705 0.103
5.4 Extra Findings
We replace the embeddings obtained by the CLIP vision encoder and the alignment module with the embeddings directly from the CLIP vision encoder without passing them through the alignment module. The results are shown in Table 3. Interestingly, we find that using the embeddings directly from the CLIP vision encoder can even have slightly better performance. However, the CLIP encoder is frozen during the training stage of LLaVA v 1.5. This suggests that the images were likely included in the CLIP model’s original training data. Our method’s high performance might also benefit from this. This suggests that we need some new benchmark datasets to better define the image MIA problem in the context of LVLMs.
6 Conclusion
In this paper, we design novel membership inference attack methods named ICIMIA against LVLMs under both white-box and black-box settings. Our approach is based on the observation that LVLMs exhibit varying sensitivity to image corruption for member and non-member images. We leverage this phenomenon by using the pair-wise similarity of the original version and its corrupted counterpart as the metric. Experimental results on existing datasets validate the effectiveness of our proposed methods.
Limitations
We observed an interesting phenomenon: many non-member images generated by DALL·E (Ramesh et al., 2022) exhibit greater robustness to corruption compared to their original counterparts (the generated image is prompted to be similar to its original member image), regardless of whether the original image is a member or non-member. Therefore, our method does not work for such a dataset where each non-member image is generated by DALL-E based on the original member image. One example is the img_dalle dataset Li et al. (2024b) where Blip (Li et al., 2023b) generates a caption for each member image and the caption is used by DALL-E to generate a non-member image for this image. We leave this as our future work.
Ethical Considerations
Our work aims to find the images that are in the training data of large vision-language models. Our proposed method, ICIMIA, can help individuals know whether their sensitive data is used to train the model. All experiments are conducted on open-source models and publicly available datasets.
AI Assistants In Writing. In this paper, we only use AI assistants for grammar checking and sentence polishing.
Appendix A Details of Datasets
Many models, such as LLaVA 1.5 and Minigpt 4 (Zhu et al., 2023), use MS COCO (Lin et al., 2014) to train the models. Therefore, Li et al. (2024b) use part of the images in MS COCO as the member images. For non-member data, Li et al. (2024b) select images uploaded to Flickr [^2] https://www.flickr.com/ after those models’ release date. The datasets are licensed under the Creative Commons Attribution 4.0 International License. The dataset statistics are summarized in Table 4.
Table 4: Dataset statistics. The datasets are constructed by Li et al. (2024b)
Name Member Data Non-member Data #Memebr Data #Non-member Data
VL-MIA/Flickr MS COCO Flickr 300 300
VL-MIA/Flickr 2K MS COCO Flickr 1000 1000
Appendix B Examples of Different Corruptions
In this section, we show some examples of different types of corruption. The examples are shown in Figure 9. The original image is from the VL-MIA/Flickr (Li et al., 2024b) dataset and is sourced from: https://www.flickr.com/photos/9750464@N02/53573626031
Figure 9: Examples of different corruption methods.
Appendix C Used Metrics
We provide details about two used metrics here:
  • AUC: Area Under the Curve (AUC) is the value of the area beneath the ROC curve. It is a widely used metric to evaluate the classification model’s performance under all possible classification thresholds (Li et al., 2024b).
  • TPR at 5% FPR: True Positive Rate at 5% False Positive Rate (TPR at 5% FPR) is another widely used metric to evaluate the performance of a classification model (Li et al., 2024b; Carlini et al., 2022a). It reflects the value of the True Positive Rate when the False Positive Rate is 5%.
Appendix D Additional Results
We conduct experiments on VL-MIA/Flickr-2K. The image similarity-based MIA results are shown in Table 5 and Table 6. We can have similar observations as in Table 1 and Table 2. The experimental results on VL-MIA/Flickr-2K also validate the effectiveness of our proposed methods.
Appendix E Potential Risks
Although our proposed ICIMIA is designed to protect data integrity, malicious users can use ICIMIA to infer whether an image is used to train a certain LVLM and then get some private information. For example, if the attacker knows that one person’s medical image is used to train an LVLM for a certain disease, the attacker can get the information that this person is likely to have this disease.
Table 5: Image MIA AUC performance of our proposed approach on VL-MIA/Flickr 2K.
Method LLaVA-1.5-7B LLaVA-1.5-13B
Ours (Image_Sim, Gaussian Blur, Kernel Size 5) 0.854 0.850
Ours (Image_Sim, Motion Blur, Kernel Size 5) 0.842 0.837
Ours (Image_Sim, JPEG Compression, Quality = 5) 0.629 0.635
Table 6: Image MIA TPR at 5% FPR performance of our proposed approach on VL-MIA/Flickr 2K.
Method LLaVA-1.5-7B LLaVA-1.5-13B
Ours (Image_Sim, Gaussian Blur, Kernel Size 5) 0.349 0.337
Ours (Image_Sim, Motion Blur, Kernel Size 5) 0.372 0.352
Ours (Image_Sim, JPEG Compression, Quality = 5) 0.088 0.080